# SH2: Self-Highlighted Hesitation Helps You Decode More Truthfully

### v3

@@ 4.2 Experimental Settings
We apply our SH2 on LLaMA-7b (Touvron et al., 2023a), LLaMA2-7b (Touvron et al., 2023b) and Mistral-7b Jiang et al. (2023). We compare it with other SOTA (state-of-the-art) methods, including Alpaca (Taori et al., 2023), ITI (Li et al., 2023b), 13b-CD (Chuang et al., 2023) and DoLa (Chuang et al., 2023). All of these baselines use LLaMA-7b as their backbone. It should be noted that ITI uses 5% samples of TruthfulQA in few-shot prompting. Additionally, LLaMA-13b is used in 13b-CD to be contrasted with the 7b model. We implement Alpaca and DoLa following the official instructions and report the evaluation results of our implementation.
@@
Since there are only about a dozen tokens in each question of TruthfulQA, the sample proportion $\eta$ of LLaMA-7b is set to be 10% and 40% for the discrimination track and the generation track respectively. And the drop-out rate $\lambda$ is set to be 0. While for HaluEval-Sum, which has about a thousand tokens in each document, $\eta$ and $\lambda$ are set to be 6% and 0.33 for LLaMA-7b. The settings of the hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Appendix A.
@@ 4.3 Main Results
The results on the three benchmarks are shown in Table 1. Our SH2 obtained significant and consistent improvements for LLaMA-7b, LLaMA2-7b and Mistral-7b. When compared with other SFT or decoding reformulation methods, our method outperforms them in most of the metrics of the benchmarks. There is no additional data or model used in our SH2. We ask models to select the hardest tokens and hesitate on them. Although LLaMA2 and Mistral have been aligned to generate truthful answers during its training procedure, they still benefit greatly from our inference-time method.
Moreover, our SH2 achieved SOTA on both the generation and the discrimination tracks of TruthfulQA. The scores of our method are also the highest or the second best on the other three discrimination tasks. It can be concluded from the table that our approach can effectively elicit factual knowledge inside LLMs. It can not only help LLMs distinguish factual contexts and hallucinated contexts, but also guide them to generate more truthful answers.
@@
4.4 Bias in HaluEval-Sum
Looking into the summarization track of HaluEval Li et al. (2023a), it can be noted from Table 1 that there is a huge difference between the values of Acc-A and Acc-H for LLaMA-7b, Alpaca and LLaMA2-7b. Acc-A is the average of the accuracies on hallucinated summaries and right summaries. Acc-H is computed by averaging the reciprocals of the accuracies and then taking the reciprocal of the average. Acc-A can be affected by extreme values easily, while Acc-H is more balanced.
The great gap between Acc-A and Acc-H denotes LLMs’ bias in hallucinated summaries and right summaries even for LLaMA2-7b. While our method effectively alleviates the problem.
Furthermore, we regard hallucinated summaries as positive labels and right summaries as negative. We also calculate precision, recall and F1 scores to evaluate the truthfulness more thoroughly. The scores are reported in Table 2. Precision denotes the percentage of real hallucinations the model determines to be hallucinated, while recall denotes the accuracy on hallucinated summaries. It can be learned from the table that LLaMA2-7b has more confidence in identifying factual summaries. However, it is still challenging for LLaMA2-7b not to be confused by hallucinated summaries, which is a potential risk for the development of LLMs.
Nevertheless, our method successfully mitigates this imbalance and benefits the general performance. It can advance the discriminant ability of models comprehensively.
@@ 5.1 Choices of Highlighted tokens
In contrast to key tokens that are the hardest to predict, we sample the easiest tokens with the highest prediction probability. Additionally, we also randomly sample the same number of tokens for comparison. We conduct experiments with LLaMA2-7b on 1000 samples of HaluEval-Sum. The effective sampling proportion $\eta^{\prime}=(1-\lambda)\cdot\eta$ ranges from 1% to 8% with the step of %1. Average scores and standard deviations are calculated for the three choices of highlighted tokens.
The accuracies on hallucinated summaries and right summaries, and overall scores (Acc-A and Acc-H) are shown in Figure 3a. When the hardest tokens are highlighted in hesitations, the LLaMA2-7b obtains better and more consistent performance than when highlighting the easiest tokens or randomly. It can be noticed that the accuracy on the hallucinated summaries is much better than accuracies of the other two.
Figure 3b illustrates the harmonic accuracy for different choices of highlighted tokens with respect to the sampling proportion $\eta$. The scores of the hardest tokens are consistently better than the other two. It is worth noting that only 1% key tokens to be highlighted in hesitations are helpful enough for language models to distinguish hallucinated contexts.
@@
Moreover, we compare precision, recall and F1 scores for the three choices of highlighted tokens in Figure 4. The disturbance brought by highlighting the easiest tokens or random tokens hurts the precision of LLMs, while the hardest tokens are beneficial. It can also be concluded from the figure that the manner of hesitations helps mitigate the bias in HaluEval-Sum.
@@ 5.2 Manners of Hesitations
In addition to underlining key tokens in the input text, we can also repeat the text or pause. Experiments are conducted on TruthfulQA with LLaMA-7b. For the repetition manner, we can repeat the questions in hesitations. While for the pausing manner, key tokens $T_{2}(X)$ are replaced by several pause words (".") in hesitations. The performance of pausing hesitations is evaluated with 3, 6, 9 and 12 pause words. MC2 scores with different numbers of pause words and the adjustment hyper-parameter $\alpha$ are shown in Figure 6. 6 pause words are used in hesitations in the following study.
The scores on the discrimination track and the generation track of the three manners are reported in Table 3. Hesitations with key tokens achieve the most significant improvement on both tracks, while pausing hesitations even hurt the generation quality of LLaMA-7b. It can be found from the table that all of the three manners help discriminate hallucinated answers. The results demonstrate that the difference brought by hesitations can elicit factual knowledge inside language models.
@@ 5.3 Effect of Contrastive Decoding
Our method does not use the original form of contrastive decoding. The difference of confidence brought by hesitations, which is the item of contrastive decoding, is traded off with the base probability. We conduct experiments on TruthfulQA with different values of the adjustment hyper-parameter $\alpha$ to study the effect of contrastive decoding (CD).
The effect of contrastive decoding with different numbers of highlighted tokens is shown in Figure 5. Since there are only about a dozen tokens in each question of TruthfulQA, the hesitation contains only one token when sampling the top 10% hardest tokens. LLaMA-7b benefits a lot from the one or two highlighted tokens.
When $\alpha$ is set to be 0, the model directly decodes with the input text followed by the hesitation. The MC2 score even drops a little compared to the vanilla LLaMA-7b. As more weight is put on the difference brought by hesitations, the model obtains a more significant improvement. Specifically, when the weight is big enough to ignore the base probability, the decoding procedure is equivalent to standard contrastive decoding without the base probability as the equation (5). It can be concluded from the figure that, the highlighted tokens make a difference in the confidence to identify factual knowledge.
@@ 6 Conclusion
In this paper, we study the bias in capturing important factual information for large language models and introduce a novel inference-time method SH2. Our method proposes to highlight the key tokens that are hard for LLMs to predict and construct hesitations with these informative tokens. To help LLMs capture critical semantic information, the decoding procedure is reformulated with the probability difference brought by hesitations. Comprehensive experiments and analysis are conducted on multiple tasks. SH2 significantly improves the truthfulness of LLMs with no need for external data or models. In the age of LLMs, our SH2 could inspire a new paradigm for faithful LLMs.
@@ Appendix A Hyper-parameter Settings
The settings of the hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Table 4. The drop-out rate $\lambda$ is set to be 0 for the discrimination track and the generation track of TruthfulQA, whose questions are of a dozen tokens. It is set to be 0.33 for FACTOR and HaluEval-Sum, because these benchmarks have longer documents.
@@ Appendix B Multiple-Choice Metrics
For the discrimination track of TruthfulQA, we use MC1, MC2 and MC3 scores to measure the truthfulness of a language model. The definitions of each score are as follows.
  • MC1: Among the set of true and false reference answers, we need to choose the best correct answer. MC1 is computed by whether the language model assigns the highest likelihood to the best correct answer over the false answers given the question.
@@
  • MC3: MC3 is computed by whether the language model assigns a higher likelihood to the correct answers over the false answers given the question.
Appendix C Constrastive Decoding
C.1 Effect of CD with Pause Words
@@
C.2 Case Study
A case study is given in Figure 7 to illustrate the effect of our self-highlighted hesitations and contrastive decoding. We calculate the probability given by LLaMA-7b. $p_{\theta}(Y|X)$ and $p_{\theta}(Y|X^{\prime})$ stand for standard decoding without and with the hesitation for each option. The difference in probability are derived by $\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)}$.
It can be learned from the figure that, with hesitations, LLaMA-7b is prone to assign higher probabilities to correct answers and lower probabilities to false answers. The model is more confident in selecting correct answers and rejecting false answers. It demonstrates that CD helps LLMs distinguish correct answers from false ones. It plays a role in separating the positive and negative answers.
Appendix D Key Tokens
D.1 Normalized Top-$\eta$ Recall
@@
D.2 Visualization of Token Probabilities

### v4

@@ 4.2 Experimental Settings
We apply our SH2 on LLaMA-7b (Touvron et al., 2023a), LLaMA2-7b (Touvron et al., 2023b) and Mistral-7b Jiang et al. (2023). We compare it with other SOTA (state-of-the-art) methods, including Alpaca (Taori et al., 2023), ITI (Li et al., 2023b), 13b-CD (Chuang et al., 2023) and DoLa (Chuang et al., 2023). All of these baselines use LLaMA-7b as their backbone. It should be noted that ITI trained a probe with the data of TruthfulQA to assist the inference of LLaMA. LLaMA-13b is used as the expert model in 13b-CD to be contrasted with the 7b model. We implement Alpaca and DoLa following the official instructions and report evaluation results of our implementation.
@@
Since there are only about a dozen tokens in each question of TruthfulQA, the sample proportion $\eta$ of LLaMA-7b is set to be 10% and 40% for the discrimination track and the generation track respectively. The drop-out rate $\lambda$ is set to be 0. For HaluEval-Sum, which has about a thousand tokens in each document, $\eta$ and $\lambda$ are set to be 6% and 0.33 for LLaMA-7b. The settings of hyper-parameters are summarized in Appendix A.
@@ 4.3 Main Results
The results on the three benchmarks are shown in Table 1. Our proposed SH2 exhibits noteworthy and consistent enhancements across LLaMA-7b, LLaMA2-7b, and Mistral-7b. SH2 outperforms other SFT or decoding reformulation techniques in the majority of metrics across these tasks. Notably, SH2 does not require any external data or model. It only asks LLMs to select the hardest tokens and hesitate on them. Even for models like LLaMA2 and Mistral, which have undergone truthfulness alignments during their training, our inference-time method can still yield substantial gains.
Moreover, our SH2 achieves SOTA on both the generation and discrimination tracks of TruthfulQA. The scores of our method are either the highest or the runner-up in the remaining three tasks. The table suggests that our approach can effectively elicit factual knowledge inside LLMs. It can not only help LLMs distinguish factual and hallucinated contexts, but also guide them to generate more truthful answers.
@@
4.4 LLMs’ Bias in HaluEval-Sum
Upon examining the summarization track of HaluEval, a significant discrepancy is observed between the Acc-A and Acc-H scores for LLaMA-7b, Alpaca, and LLaMA2-7b, as shown in Table 1. Acc-A, which represents the average accuracy on hallucinated and right summaries, is susceptible to extreme values. Conversely, Acc-H, calculated by averaging the reciprocals of accuracies and then taking the reciprocal of the average, provides a more balanced assessment.
The value discrepancy between Acc-A and Acc-H denotes LLMs’ bias towards hallucinated and right summaries. However, our method has been shown to effectively address this issue.
We evaluate the truthfulness more thoroughly by considering hallucinated summaries as positive labels and right summaries as negative, and calculating precision, recall and F1 scores as reported in Table 2. Precision denotes the percentage of real hallucinations the model determines to be hallucinated, while recall denotes the accuracy on hallucinated summaries. The high precision and low recall indicate that LLaMA2-7b exhibits high confidence in identifying factual summaries. Yet, it remains a considerable challenge for LLaMA2-7b to distinguish hallucinated summaries, posing a potential risk for the development of LLMs.
Nevertheless, our method has proven effective in reducing this discrepancy and enhancing overall performance. It could advance the discriminant ability of models comprehensively.
@@ 5 Analysis
We conduct further studies regarding choices of highlighted tokens, manners of hesitations, and the effect of contrastive decoding in this section.
@@ 5.1 Choices of Highlighted tokens
In contrast to key tokens that are the hardest to predict, we sample the easiest tokens with the highest prediction probability. Additionally, we also sample the same number of tokens randomly for comparison. We conduct experiments with LLaMA2-7b on 1000 samples of HaluEval-Sum. The effective sampling proportion $\eta^{\prime}=(1-\lambda)\cdot\eta$ ranges from 1% to 8% with the step of %1. Average scores and standard deviations are calculated for the three choices of highlighted tokens.
The accuracies on hallucinated summaries and right summaries, and overall scores (Acc-A and Acc-H) are shown in Figure 3a. When the hardest tokens are highlighted in hesitations, the LLaMA2-7b obtains superior and more consistent performance compared to highlighting the easiest tokens or randomly, particularly in the case of hallucinated summaries.
Figure 3b illustrates the harmonic accuracy for different choices of highlighted tokens with respect to the sampling proportion $\eta^{\prime}$. The scores by selecting the hardest tokens remain consistently higher than those of the other two choices. It is noteworthy that highlighting only 1% tokens with the lowest prediction probabilities during hesitations is sufficient enough for the model to distinguish hallucinated contexts.
@@
Moreover, we compare precision, recall and F1 scores for the three choices of highlighted tokens in Figure 4. The results indicate that highlighting the easiest tokens or random tokens negatively impacts the precision of LLMs. However, highlighting the hardest tokens is beneficial. It effectively mitigates LLMs’ bias in HaluEval-Sum.
@@ 5.2 Manners of Hesitations
In addition to underlining key tokens in the input text, we can also repeat the text or pause. Experiments are conducted on TruthfulQA with LLaMA-7b. For the repetition manner, we can simply repeat the question as hesitations. As for the pausing manner, several pause words (".") are appended to the question as hesitations. The performance of pausing hesitations is evaluated with 3, 6, 9 and 12 pause words. 6 pause words have the best performance and are used in hesitations in the following study. Scores with different numbers of pause words and the adjustment hyper-parameter $\alpha$ are shown in Appendix D.2.
The results on the discrimination track and the generation track of the three manners are reported in Table 3. Hesitations with key tokens achieve the highest scores on both tracks, while pausing hesitations even hurt the generation quality of LLaMA-7b. Furthermore, the improvements of MC scores demonstrate that all of the three manners are effective in distinguishing hallucinated answers. It suggests that the difference brought by hesitations could elicit factual knowledge inside LLMs.
@@ 5.3 Effect of Contrastive Decoding
Our method does not adhere to the original form of contrastive decoding, which contrasts the probabilities of two sources. As elaborated in Section 3.4, we use the parameter $\alpha$ to balance the contrastive term with the base probability. To investigate the effect of contrastive decoding, we conduct experiments on TruthfulQA with varying $\alpha$.
Figure 5 depicts the effect of contrastive decoding when utilizing different numbers of highlighted tokens. Given that each question in TruthfulQA comprises approximately a dozen tokens, the hesitation typically contains a single token when sampling the top 10% hardest tokens. It can be observed that LLaMA-7b significantly benefits from the single highlighted token.
When $\alpha$ is set to 0, the model directly decodes with the input text followed by the hesitation. The MC2 score slightly decreases compared to the baseline. However, as more weight is put on the contrastive term, the model obtains more pronounced improvements. Specifically, when $\alpha$ is large enough to overshadow the base probability, the decoding procedure is equivalent to standard contrastive decoding without the base probability as the equation (5). The figure suggests that the contrastive term makes a difference in identifying factual knowledge.
@@ 6 Conclusion
In this paper, we delve into the challenge of large language models when capturing important factual information. To address this, we introduce a novel inference-time method, SH2. Our method proposes to highlight the key tokens that are hard for LLMs to predict and construct hesitations with these informative tokens. By reformulating the decoding procedure with the probability differences brought by hesitations, we enable LLMs to discern factual content more effectively. Through extensive experiments and analysis across multiple tasks, SH2 demonstrates a significant enhancement in the truthfulness of LLMs, all achieved without reliance on external data or models.
@@
Acknowledgement
This work was sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402) and National Natural Science Foundation of China (NSFC) grant (No.62106143).
@@ Appendix A Hyper-parameter Settings
The settings of hyper-parameters $\eta$, $\lambda$ and $\alpha$ are summarized in Table 5. The drop-out rate $\lambda$ is set to be 0 for the discrimination track and the generation track of TruthfulQA, whose questions are of a dozen tokens. It is set to be 0.33 for FACTOR and HaluEval-Sum because these benchmarks have longer documents.
Models MC1 MC2 MC3 Wiki-Factor
LLaMA-7b 23.62 41.21 19.33 58.55
+SH2 27.91 55.63 29.73 63.06
LLaMA-13b 28.27 43.32 20.85 62.73
+SH2 31.09 54.31 27.16 65.80
LLaMA2-7b 28.40 43.39 20.53 58.65
+SH2 33.90 57.07 29.79 64.09
LLaMA2-13b 29.01 44.27 20.71 64.03
+SH2 37.58 62.82 34.15 66.00
Table 4: Performance of 7b and 13b models.
Parameters TruthfulQA FACTOR HaluEval
Discrimination Generation Wiki News Summarization
$\lambda$ 0 0 0.33 0.33 0.33
LLaMA-7b $\eta$ 10% 40% 24% 12% 6%
$\alpha$ 6 3.7 0 0.1 1.6
LLaMA2-7b $\eta$ 20% 30% 24% 18% 3%
$\alpha$ 27 3.4 0 0 1.6
Mistral-7b $\eta$ 25% - 18% 12% 4.5%
$\alpha$ 9 - 0 0.1 2.2
Table 5: Settings of Hyper-parameters on each task.
@@ Appendix B Multiple-Choice Metrics
For the discrimination track of TruthfulQA, we use MC1, MC2 and MC3 scores to measure the truthfulness of a language model. The definitions of each metric are as follows.
  • MC1: Among the set of true and false reference answers, the language model needs to choose the best correct answer. MC1 is computed by whether the model assigns the highest likelihood to the best correct answer over false answers given the question.
@@
  • MC3: MC3 is computed by whether the model assigns a higher likelihood to correct answers over false answers given the question.
Appendix C Extending to larger model sizes
We the discrimination track of use TruthfulQA (MC1, MC2, and MC3) and Wiki-Factor to test the performance of our method when applied to the larger models (LLaMA-13b and LLaMA2-13b) on a 40G-A100. The results are shown in Table 4. It demonstrates that our method is still effective on the larger models.
@@ Appendix C Extending to larger model sizes
Appendix D Constrastive Decoding
D.1 Case Study
A case study is given in Figure 6 to illustrate the effect of our self-highlighted hesitation and contrastive decoding. We calculate the probability given by LLaMA-7b. $p_{\theta}(Y|X)$ and $p_{\theta}(Y|X^{\prime})$ stand for standard decoding without and with the hesitation for each option. The differences in probability are derived by $\frac{p_{\theta}(Y|X^{\prime})}{p_{\theta}(Y|X)}$.
It can be learned from the figure that, with hesitations, LLaMA-7b is prone to assign higher or similar probabilities to correct answers and lower probabilities to false answers. The model is more confident in selecting correct answers and rejecting false answers. It demonstrates that CD helps LLMs distinguish correct answers from false ones. It plays a role in separating the positive and negative answers.
However, contrastive decoding might exhibit limitations when dealing with high-confidence answers like (b), whose probability change brought by hesitations is marginal. In consequence, the base probability $p_{\theta}(Y|X^{\prime})$ will predominate $p_{\mathbf{H}}(Y|X)$ to help LLMs make predictions as elaborated in Section 3.4.
D.2 Effect of CD with Pause Words
@@
Appendix E Highlighting Words of Certain POS Tags
As introduced in Section 3.2, compared with function words like IN, DT, and CC, content words like NN, NNP, and JJ are more difficult for LLMs to predict and more concentrated in the hardest part. We conducted experiments to study whether words of certain POS tags highlighted in hesitations can improve the truthfulness of LLMs.
We count the number of different pos tags in 1000 documents of HaluEval-Sum. The top five tags are NN, IN, NNP, DT, and JJ. Specifically, NN, NNP, and JJ are content words and have high normalized top-$\eta$ recall as shown in Figure 2. On the other hand, IN and DT are function words whose normalized recall scores are much lower. We sample the hardest tokens with each of these five tags. We conduct experiments with LLaMA2-7b on 1000 samples of HaluEval-Sum. The effective sampling proportion $\eta^{\prime}=(1-\lambda)\cdot\eta$ ranges from 1% to 5% with the step of %1. Average scores and standard deviations are calculated for different choices of POS tags.
Figure 8: Arithmetic-mean accuracy (Acc-A) and harmonic-mean accuracy (Acc-H) of LLaMA2-7b when highlighting words of certain POS tags. The errorbar denotes standard deviations. The effective sampling proportion $\eta^{\prime}$ ranges from 1% to 5%. The errorbar denotes standard deviations. "SH2" stands for the standard SH2 with no requirement on POS tags. We use dark colors for function words (IN and DT), and light color for content words (NN, NNP and JJ)
The scores of each POS tag are presented in Figure 8. The figure reveals that when highlighting words belonging to certain POS tags, the potential for improvement is somewhat restricted. Although NNP contributes the most among the five, it still exhibits large variance, indicating that its performance is not uniformly consistent. This inconsistency is likely attributed to the varying distribution of POS tags across different documents. Words of different POS tags could carry different information. By not overly relying on any single type of POS tag, our SH2 is more balanced and robust.
Appendix F Reasoning Ability and Inference Overhead
F.1 Reasoning on StrategyQA
Besides truthfulness, we conduct experiments on StrategyQA Geva et al. (2021) to evaluate whether our method can improve the reasoning ability of LLMs. StrategyQA contains 2290 questions requiring a multi-hop strategy for answers. The COT prompts we used are from Wei et al. (2022). We test the accuracies with original questions (w/o. COT) and with COT-prompted questions (w. COT). The results are recorded in Table 6.
It can be observed from the table that our SH2 enhances the reasoning ability of LLMs. The SFT method, Alpaca, achieves the best scores when no demonstrations of COT are provided. This is probably because it has been fine-tuned on 52K instruction-following data. It has developed the explicit ability of reasoning and does not benefit much from COT. Designed by humans, COT can provide additional information and teach LLMs to reason. With no additional training or external data, our method is still comparable. It even achieves superior performance when integrated with COT.
Models w/o. COT w. COT
LLaMA-7b 51.22 60.48
+Alpaca 60.70 61.62
+SH2 58.73 61.05
LLaMA2-7b 52.62 60.74
+SH2 55.15 61.88
Table 6: Accuracies (%) on StrategyQA. 6 demonstrations are used in our experiments.
F.2 Inference Overhead
Models additional data or model accuracy (%) memory (GB) time (s)
LLaMA-7b (naive decoding) None 51.22 13.69 0.32
+13b-CD Llama-13B 55.11 39.02 0.66
+DoLa None 50.79 13.81 0.34
+DoLa+COT COT data 64.15 14.30 1.89
+SH2 None 58.73 13.95 1.04
Table 7: Inference overhead statistics of SH2 and and other decoding strategies.
Without additional training or interaction with external tools in our SH2, we give an analysis of the inference overhead. We compare our method with naive decoding, the standard CD which contrasts LLaMA-13b with the 7b model, and the sota CD method DoLa which only uses the base model like our method. We use StrategyQA to evaluate the computation cost and performance. All the experiments are conducted in the same environment. Table 7 reports the average accuracy, inference time, and memory cost on these 2290 samples.
13b-CD loads the 7b and 13b models simultaneously on the GPU and sacrifices much more memory. While DoLa only works with the help of COT, our method outperforms the other two CD methods and can also be integrated with COT as in Table 6. The inference overhead of our method mainly comes from two calls of the forward function caused by contrastive decoding. It can be regarded as a tradeoff between performance and computation cost.
Appendix G Key Tokens
G.1 Normalized Top-$\eta$ Recall
@@
G.2 Visualization of Token Probabilities